OpenStudy (ohohaye):

how do I evaluate arccsc root2? Please help me

OpenStudy (freckles):

\[arccsc(\sqrt{2})=y \\ \sqrt{2}=\csc(y) \\ \sqrt{2}=\frac{1}{\sin(y)} \\ \frac{1}{\sqrt{2}}=\sin(y) \\ \arcsin(\frac{1}{\sqrt{2}}) =y \\ \text{ can you say what } \arcsin(\frac{1}{\sqrt{2}}) \text{ equals ? }\]

OpenStudy (phi):

notice another way to write \[ \frac{1}{\sqrt{2}} = \frac{\sqrt{2} }{2} \] you want to find the angle x , when \[ \sin x = \frac{\sqrt{2} }{2} \] this is one of the angles you should memorize

OpenStudy (ohohaye):

The thing is, I'm supposed to figure out how to put that in to something like pi/3, how would I do that?

OpenStudy (jdoe0001):

hmm look at your Unit Circle :) or use a calculator for that matter, that'd work as well

OpenStudy (ohohaye):

Do I need a graphing calculator for this?

OpenStudy (jdoe0001):

hmmm hold the mayo did you mean btw \(\Large cos^{-1}(\sqrt{2})\) ?

OpenStudy (ohohaye):

No the problem just said to evaluate arccsc ( 1/ \sqrt2)

OpenStudy (mathmath333):

graphing calculator is not necessary

OpenStudy (jdoe0001):

hmm 1/? where did that come from

OpenStudy (mathmath333):

u may look at frekel's method at the top , or just study the unit circle as stated

OpenStudy (jdoe0001):

well... still , very simple as freckles showed above that angle unit, would be in your Unit Circle

OpenStudy (ohohaye):

There isn't a circle stated with the question, what I said is all that there was

OpenStudy (mathmath333):

unit circle is the basic topic to know in trignometry. https://www.mathsisfun.com/geometry/unit-circle.html

OpenStudy (ohohaye):

 so something like this?

OpenStudy (mathmath333):

it looks like a pizza

OpenStudy (ohohaye):

Well you get what I'm going for right?

OpenStudy (mathmath333):



OpenStudy (mathmath333):

it s not upto 9pi its upto 0 to 360\((2\pi)\)

OpenStudy (mathmath333):

in unit circle u need to learn the basic values of angles of \(0,30,45,90\) degrees of \(\sin,\tan \cos\) etc.

OpenStudy (jdoe0001):

well...\(\bf cos^{-1}\left( \cfrac{1}{\sqrt{2}} \right)\qquad thus \\ \quad \\ \cfrac{1}{\sqrt{2}} \cdot \cfrac{\sqrt{2}}{\sqrt{2}}\implies \cfrac{1\cdot \sqrt{2}}{\sqrt{2}\cdot \sqrt{2}}\implies \cfrac{\sqrt{2}}{\sqrt{2^2}}\implies \cfrac{\sqrt{2}}{2}\qquad thus \\ \quad \\ cos^{-1}\left( \cfrac{1}{\sqrt{2}} \right)\iff cos^{-1}\left( \cfrac{\sqrt{2}}{2} \right)\impliedby \begin{array}{llll} \textit{the angle whose cosine}\\ is\quad \cfrac{\sqrt{2}}{2} \end{array}\)